solve. approximate irrational roots to the nearest tenth.
The difference of the square of a number and the number is 12. find the number.

Answer by fractalier(2101)

Let the number be x...then we have
x^2 - x = 12
x^2 - x - 12 = 0
(x - 4)(x + 3) = 0
x = 4 or x = -3